# HEART: Achieving Timely Multi-Model Training for Vehicle-Edge-Cloud-Integrated Hierarchical Federated Learning

### v1

@@ Abstract
The rapid growth of AI-enabled Internet of Vehicles (IoV) calls for efficient machine learning (ML) solutions that can handle high vehicular mobility and decentralized data. This has motivated the emergence of Hierarchical Federated Learning over vehicle-edge-cloud architectures (VEC-HFL). Nevertheless, one aspect which is underexplored in the literature on VEC-HFL is that vehicles often need to execute multiple ML tasks simultaneously, where this multi-model training environment introduces crucial challenges. First, improper aggregation rules can lead to model obsolescence and prolonged training times. Second, vehicular mobility may result in inefficient data utilization by preventing the vehicles from returning their models to the network edge. Third, achieving a balanced resource allocation across diverse tasks becomes of paramount importance as it majorly affects the effectiveness of collaborative training. We take one of the first steps towards addressing these challenges via proposing a framework for multi-model training in dynamic VEC-HFL with the goal of minimizing global training latency while ensuring balanced training across various tasks—a problem that turns out to be NP-hard. To facilitate timely model training, we introduce a hybrid synchronous-asynchronous aggregation rule. Building on this, we present a novel method called Hybrid Evolutionary And gReedy allocaTion (HEART). The framework operates in two stages: first, it achieves balanced task scheduling through a hybrid heuristic approach that combines improved Particle Swarm Optimization (PSO) and Genetic Algorithms (GA); second, it employs a low-complexity greedy algorithm to determine the training priority of assigned tasks on vehicles. Experiments on real-world datasets demonstrate the superiority of HEART over existing methods.
Index Terms:
Hierarchical federated learning, Internet of Vehicles, Multi-model training, Distributed machine learning.
@@ I Introduction
The rapid advancement of Artificial Intelligence (AI), communication and computing technologies, as well as evolving vehicle design requirements have led to the integration of Machine Learning (ML)-driven applications within the Internet of Vehicles (IoV) [1, 2], aiming to enhance vehicular services and driving experiences. However, the distributed nature of data generated across the IoV and the high mobility of vehicles introduce significant challenges to the efficient utilization of IoV data for ML training purposes. This is due to the fact that conventional ML training methods often rely on centralized procedures that entail transferring data from all data-generating units (e.g., vehicles) to a single location (e.g., a cloud server). This approach results in high latency, increased network traffic congestion, and raises privacy concerns due to the exposure of raw data. Federated Learning (FL) has emerged as an innovative distributed ML framework that addresses these issues by pushing model training to the data-generating units themselves, thereby preserving data privacy and reducing communication overhead [3]. Nevertheless, in dynamic vehicular environments, FL still suffers from low communication efficiency and frequent communication failures [4, 5, 6] due to the high mobility of vehicles and unstable network connections. To overcome these challenges, Hierarchical Federated Learning (HFL) has been proposed as a multi-layer extension of FL [7]. HFL introduces intermediate aggregators (e.g., edge servers) between the vehicles and the cloud server, helping to balance the trade-off between communication efficiency and computational overhead. This hierarchical approach enhances the robustness and efficiency of the learning process, making HFL a promising solution for efficient ML training and execution in IoV systems.
I-A Integration of FL and HFL into the IoV Ecosystem
@@
Recent integrations of conventional FL and HFL over IoV systems predominantly emphasize single-task/model FL [8, 9, 10], where all vehicles are engaged in training a single ML model. However, in real-world vehicular networks, vehicles often need to execute multiple ML tasks simultaneously, such as object detection for autonomous driving, traffic prediction for route optimization, and driver behavior analysis for safety enhancements. This necessitates the use of multi-model training approaches that can handle diverse ML models concurrently. Despite its promising potential, there exists very limited literature addressing multi-model FL over wireless networks [11, 12], and these works often rely on conventional FL training methods, which face significant communication overhead. Consequently, reducing communication costs while accelerating the overall multi-model training process is a primary concern that demands critical attention for both IoV systems and wireless networks more broadly.
@@
Some early efforts have been dedicated to address these concerns, focusing on the conventional FL architecture. For instance, recent studies have considered the unique characteristics of diverse ML models by leveraging vehicle-side resources to reduce the required number of global iterations, thereby ensuring timely and efficient model training [13, 14]. In these approaches, local models are collectively transmitted only once when their training is finalized on the vehicles, effectively minimizing the frequency of communication with the global aggregator, e.g., a cloud server [6]. Additionally, some research highlights the importance of selecting appropriate clients for training and effectively aggregating results from various tasks in FL[15]. While these contributions have advanced the new field, they often overlook the impact of balanced scheduling of ML tasks on global model convergence when handling multiple tasks with a HFL over dynamic IoV systems. In such systems, risks may arise due to model obsolescence if the trained models are not promptly aggregated upon completion [5]. Further, improper prioritization of task training on vehicles can lead to prolonged waiting times for model aggregation, further exacerbating potential model obsolescence.
@@ I-B Motivation and Overview of Our Methodology
This work is among the first in the literature to address the multi-model training within VEC-HFL. In our problem of interest, vehicles serve as training clients, multiple edge servers (ESs) act as middle-layer aggregators, and a cloud server (CS) functions as the global aggregator. We first note that a key challenge in facilitating multi-model training is mitigating model obsolescence risks while enhancing the utilization of distributed and dynamic training data from vehicles. In particular, we observe that traditional model aggregation methods in FL and HFL are either synchronous or asynchronous. Synchronous aggregation waits for all clients to upload their models to maintain consistency, but this can introduce significant delays and increase the risk of model obsolescence. Conversely, asynchronous aggregation initiates the aggregation process as soon as a subset of clients uploads their models, effectively reducing latency. However, this method may suffer from model instability and struggle to fully utilize data from all participating clients, potentially impacting the convergence speed of the global model. To address these challenges, we introduce a hybrid synchronous-asynchronous aggregation rule for VEC-HFL. This rule operates on a per-task basis: at the edge layer, each ES performs an aggregation only after receiving all local models of a specific task from vehicles (synchronous approach), while at the cloud layer, the CS conducts a global aggregation after obtaining a subset of the corresponding ES models (asynchronous approach). This hybrid approach leverages the advantages of both synchronous and asynchronous methods. By synchronously aggregating at the edge layer, we ensure model consistency and fully utilize the local training data from all participating vehicles, enhancing model accuracy and convergence speed. Simultaneously, the asynchronous aggregation at the cloud layer reduces overall latency by not waiting for all ESs, which can have different delays of model aggregation based on the distribution of vehicles covered by them and their channel conditions, to report their models, thereby mitigating delays and the risk of model obsolescence. This combination provides a natural solution that balances the trade-offs between timeliness and model quality, making it particularly suitable for dynamic IoV environments where both factors are critical.
@@
Building on this novel hybrid aggregation rule, we then target minimizing the global time/latency costs while ensuring balanced training across diverse ML tasks over resource-limited vehicles. We formulate this problem as a min-max optimization that turns out to be NP-hard, making the obtaining of its solution challenging in dynamic IoV systems. To tackle this problem, we propose a stagewise solution called the Hybrid Evolutionary And gReedy allocaTion method (HEART). In Stage 1 of HEART, we optimize the assignment between model training tasks and heterogeneous resource-limited moving vehicles. To this end, leveraging the strong global search capabilities and computational efficiency of Particle Swarm Optimization (PSO) and Genetic Algorithms (GA), we develop an efficient combination of improved PSO and GA to achieve or closely approximate the optimal and balanced task assignment solution. In Stage 2 of HEART, we coordinate the training sequence/rank of assigned tasks on each selected vehicle to accelerate decision-making by designing a greedy-based task ranking optimization method.
@@ I-C Summary of Contributions
$\bullet$ Hybrid Synchronous-Asynchronous Aggregation Rule for Multi-Model Training in VEC-HFL: We introduce a novel aggregation rule tailored for VEC-HFL. This hybrid approach performs synchronous aggregation at the ES and asynchronous aggregation at the CS. By doing so, it effectively leverages the rich data generated by vehicles while significantly mitigating the adverse effects of model obsolescence, thereby enhancing the overall timeliness of the training process across ML tasks.
@@ II Background and Preliminaries
The VEC-HFL architecture of our interest (see Fig. 1) comprises one CS as the global aggregator, multiple ESs as middle-layer aggregators, and multiple vehicles as model training clients moving between ESs. We collect the ESs via the set =$\mathcal{M}=\{1,...,M\}$, where each ES $m\in\mathcal{M}$ has a certain communication coverage. Also, we denote the set of vehicles in the IoV system by $\mathcal{N}=\{1,...,N\}$, where each vehicle $n\in\mathcal{N}$ is assumed to be located inside the considered IoV region during the training process, moving between the ESs. Moreover, we denote the set of all ML tasks in the system by $\mathcal{J}=\{1,...,J\}$, which are expected to be scheduled and trained/executed over moving vehicles.
@@
[…]
@@ III-A Modeling the Time Overhead of Local and Edge Iterations
The non-task-training time primarily consists of (i) the local model upload time and (ii) any idle period from the completion of one task’s local iteration to the start of the next [^1] We have disregarded the communication delay of the downlink between ESs-to-vehicles and CS-to-ESs, due to the strong broadcasting capabilities of ESs and CS – similar to [14, 16].. However, modeling this time needs careful considerations because, during the $g^{\text{th}}$ global iteration, vehicles immediately transmit their local models to the associated ES once they finish training, and thus two scenarios may arise:
@@
where $r_{m,n;[g,k]}^{V2E,\left\langle j\right\rangle}$ represents the vehicle-to-edge data rate of vehicle $n$ under the coverage of ES $m$ during the $k^{\text{th}}$ edge iteration, which is related to the distance between the vehicle and ES and the channel conditions [17]. Besides, the size of the local model (in bits) is denoted by $I_{m,n}^{V2E,\left\langle j\right\rangle}$.
Let $t^{inact,\langle j\rangle}_{m,n;[g,k]}$ denote the inactive time of the $n^{\text{th}}$ vehicle covered by ES $m$ after completing its local iteration at edge iteration $k$ of global aggregation $g$. It can be observed that this inactive time is a function of the completion time of edge iteration $k-1$ within global aggregation $g$ for a task $j$, which we denote it by $t_{m,n;[g,k-1]}^{end,\langle j\rangle}$. Mathematically, we obtain this inactive time as follows:
@@
$t_{m,n;[g,k]}^{end,\langle j\rangle}=\sum_{j^{\prime}:~{}\mathcal{I}(\lambda_{m,n;[g]}^{\langle j^{\prime}\rangle})\leq\mathcal{I}(\lambda_{m,n;[g]}^{\langle j\rangle})}t_{m,n;[g,k]}^{\langle j^{\prime}\rangle}\overline{x}_{m,n;[g]}^{\left\langle j^{\prime}\right\rangle},j\in\mathcal{J},$ (12)
@@ IV-A Key Constraint Design
$\frac{N}{J}-\xi_{1}\leq\sum_{m\in\mathcal{M}_{g}^{\left\langle j\right\rangle}}\sum_{n\in\mathcal{N}_{m;[g,k]}^{\left\langle j\right\rangle}}{\overline{x}_{m,n;[g]}^{\left\langle j\right\rangle}}\leq\frac{N}{J}+\xi_{2},\forall j\in\mathcal{J}.$ (17)
@@ IV-C Proposed Stagewise Scheduling through Hybrid Evolutionary And gReedy allocaTion (HEART)
Due to the integer/discrete nature of the optimization variables, $\bm{\mathcal{P}}$ belongs to a sub-category of integer programming problems which are known to be NP-hard. To tackle this problem, we propose a stagewise scheduling method called HEART, designed to achieve a near-optimal solution in a timely manner.
@@
Stage 1: Task Scheduling via a Hybrid Heuristic Approach
Heuristic algorithms such as PSO and GA are widely used due to their exceptional performance in scheduling and resource allocation problems [18, 19]. However, the complexity inherent in these problems, particularly the one considered in Stage 1 of this paper, often causes these algorithms to get trapped in local optima when they are used in isolation. To overcome this challenge, we propose a hybrid optimization algorithm that leverages the complementary strengths of both PSO and GA. In conventional PSO, particles tend to cluster prematurely, which weakens the algorithm’s global search capability [20, 21]. Meanwhile, GA requires careful parameter tuning and frequently suffers from premature convergence [22, 23]. By integrating these two methods, we strive to achieve solutions that are globally optimal—or at least near-optimal—and avoid the pitfalls commonly associated with using a single method.
@@
S1.1 Overview and sketch of the PSO process: In our proposed task scheduling framework, a conventional PSO algorithm would search for optimal scheduling solutions across all vehicles at once [18, 20]. In particular, conventional PSO, inspired by the social behavior of bird flocks or fish schools, operates by having particles (representing solutions) move through the solution space. Each particle adjusts its position based on its own experience and that of its neighbors, guided by a fitness function that evaluates solution quality. The algorithm iteratively refines particle positions to converge on an optimal or near-optimal solution. While this global approach captures overall system objectives, it significantly increases the algorithm’s complexity and makes it difficult to tailor the scheduling solution for each individual vehicle in our setting of interest. Conversely, assigning tasks on a per-vehicle basis risks neglecting the overarching need for balanced global task allocation. A particular challenge arises because assigning more tasks can provide richer training data—beneficial for building a high performance global model—but since different tasks require different training times, within a limited dwell time, vehicles may often end up choosing tasks with shorter training durations to fit in as many tasks as possible when task assignments are made on a per-vehicle basis. As a result, tasks that take longer to train will be assigned less frequently, slowing down the progression of the global model accuracy and increasing thr total training time. In the following, we detail our proposed enhanced PSO.
@@
$f_{m,n;[g]}=\begin{cases}-\infty&\text{if it can not meet (16)}\\ \sum_{j=1}^{J}\overline{x}_{m,n;[g]}^{\left\langle j\right\rangle}-(\xi_{3}\sum_{j=1}^{J}|\psi_{j}\\ ~{}~{}-\chi|-\sum_{j=1}^{J}\rho_{j}|\psi_{j}-\chi|)&\text{otherwise}\end{cases}.$ (19)
@@
$\overline{x}_{m,n;[g]}^{p,\left\langle j\right\rangle}(\tau)$ =$=\begin{cases}1&\text{if}~{}~{}e_{1}\leq\Phi^{p,\left\langle j\right\rangle}_{m,n;[g]}(\tau)\\ 0&\text{otherwise}\end{cases},$ (23)
@@
$\bullet$ Crossover enhancement. During the task scheduling phase for each vehicle, we sequentially select two adjacent particles, $p$ and $p+1$, correspond to task scheduling $\overline{x}_{m,n;[g]}^{{p,\left\langle j\right\rangle}}(\tau)$, $\overline{x}_{m,n;[g]}^{{p+1,\left\langle j\right\rangle}}(\tau)$. We then apply crossover and mutation on these allocations repeatedly. Specifically, for crossover, $p$ and $p+1$ serve as parent solutions, and we define a genetic length $J$ along with a randomly chosen crossover point $r$ such that $1<r<J$. Exchanging the segments of the parent allocations at the crossover point creates two offspring, $c_{1}$ and $c_{2}$. Thus, new task scheduling schemes are generated, which expand the exploration space and reduce the likelihood of trapping the algorithm in suboptimal solutions. We define all task scheduling for $p$, $p+1$ as a vector $X^{p}_{m,n;[g]}(\tau)$ and $X^{p+1}_{m,n;[g]}(\tau)$ (e.g., for $p$ during the $\tau^{\text{th}}$ iteration, if the particle implies the scheduling of tasks 1, 3, and 4 out of the four available tasks in the system, we will have the vector $X^{p}_{m,n;[g]}(\tau)=[1,0,1,1]$). To better describe this crossover process, we use $[i:j]$ to represent the elements from $i$ to $j-1$ in the vector, e.g., considering $X^{p}_{m,n;[g]}(\tau)=[1,0,1,1]$, we have $X^{p}_{m,n;[g]}(\tau)[0:2]=1,0$. We then obtain the offspring inherit allocations from the parents according to the following crossover rules:

### v2

@@ Abstract
The rapid growth of AI-enabled Internet of Vehicles (IoV) calls for efficient machine learning (ML) solutions that can handle high vehicular mobility and decentralized data. This has motivated the emergence of Hierarchical Federated Learning over vehicle-edge-cloud architectures (VEC-HFL). Nevertheless, one aspect which is underexplored in the literature on VEC-HFL is that vehicles often need to execute multiple ML tasks simultaneously, where this multi-model training environment introduces crucial challenges. First, improper aggregation rules can lead to model obsolescence and prolonged training times. Second, vehicular mobility may result in inefficient data utilization by preventing the vehicles from returning their models to the network edge. Third, achieving a balanced resource allocation across diverse tasks becomes of paramount importance as it majorly affects the effectiveness of collaborative training. We take one of the first steps towards addressing these challenges via proposing a framework for multi-model training in dynamic VEC-HFL with the goal of minimizing global training latency while ensuring balanced training across various tasks—a problem that turns out to be NP-hard. To facilitate timely model training, we introduce a hybrid synchronous-asynchronous aggregation rule. Building on this, we present a novel method called Hybrid Evolutionary And gReedy allocaTion (HEART). The framework operates in two stages: first, it achieves balanced task scheduling through a hybrid heuristic approach that combines improved Particle Swarm Optimization (PSO) and Genetic Algorithms (GA); second, it employs a low-complexity greedy algorithm to determine the training priority of assigned tasks on vehicles. Experiments on real-world datasets demonstrate the superiority of HEART over existing method
@@ I Introduction
The rapid advancement of Artificial Intelligence (AI), communication and computing technologies, as well as evolving vehicle design requirements have led to the integration of Machine Learning (ML)-driven applications within the Internet of Vehicles (IoV) [SplitLearning, JointClient], aiming to enhance vehicular services and driving experiences. However, the distributed nature of data generated across the IoV and the high mobility of vehicles introduce significant challenges to the efficient utilization of IoV data for ML training purposes. This is due to the fact that conventional ML training methods often rely on centralized procedures that entail data transfer from all data-generating units (e.g., vehicles) to a single location (e.g., a cloud server). This approach results in high latency, increased network traffic congestion, and growing privacy concerns due to the exposure of raw data. Consequently, Federated Learning (FL) has emerged as an innovative distributed ML framework that addresses these issues by pushing model training to the data-generating units themselves, thereby preserving data privacy and reducing communication overhead [AdaptiveTraining, GHPFL]. Nevertheless, in dynamic vehicular environments, FL still suffers from low communication efficiency and frequent communication failures [HierFedML, TranWGAN, DeviceS] due to the high mobility of vehicles and unstable network connections. To overcome these challenges, Hierarchical Federated Learning (HFL) has been proposed as a multi-layer extension of FL [11202380, ConvergenceL, WirelessP]. HFL introduces intermediate aggregators (e.g., edge servers) between the vehicles and the cloud server, helping to balance the trade-off between communication efficiency and computational overhead. This hierarchical approach enhances the robustness and efficiency of the learning process, making HFL a promising solution for efficient ML training and execution in IoV systems.
I-A Investigations on the Integration of FL and HFL within the IoV Ecosystem
@@
Recent investigations of conventional FL and HFL over IoV systems predominantly emphasize single-task/model FL [MOBFL, DistributedP, EfficientVehicle], where all vehicles are engaged in training a single ML model. However, in real-world vehicular networks, vehicles often need to execute multiple ML tasks simultaneously, such as object detection for autonomous driving, traffic prediction for route optimization, and driver behavior analysis for safety enhancements. This necessitates the use of multi-model training approaches that can handle diverse ML models concurrently. Despite its promising potential, there exists very limited literature addressing multi-model FL over wireless networks [FedICT, Asynchronous, CommunicationEfficient], and these works often rely on conventional FL training methods, which face significant communication overhead. Consequently, reducing communication costs while accelerating the overall multi-model training process is a primary concern that demands critical attention for both IoV systems and wireless networks more broadly.
@@
Some early efforts have been dedicated to address these issues, while leveraging the conventional FL architecture. For instance, recent studies have considered the unique characteristics of diverse ML models by effectively utilizing vehicle-side resources to reduce the required number of global iterations, thereby ensuring accelerated model training [MatchingGame]. In these approaches, local models are collectively transmitted only once when their training is finalized on the vehicles, effectively minimizing the frequency of communication with the global aggregator, e.g., a cloud server [AdaptiveP]. Additionally, [MASS] highlighted the importance of selecting appropriate clients for training and effectively aggregating results from various tasks in FL, [TowardsOP] proposed a loss based sampling method that can minimize the training variance during the fl process to the greatest extent possible and [ManyTask] introduced a task similarity metric based on client vector alignment, reducing communication overhead. While these contributions have advanced the new field, they often overlook the impact of balanced scheduling of ML tasks on global model convergence when handling multiple tasks with a HFL over dynamic IoV systems. To address this issue, [tang2023fedml, TaskSelection, MobilityBased] focus on solving balanced multi-task allocation problem while reducing the global train time cost in FL. However, in such systems, risks may arise due to model obsolescence if synchronous aggregation mechanism, dynamic nature of vehicles, and the trained models are not promptly aggregated upon completion [HiFlash, FedFetch, OptimizingM]. Further, improper prioritization of task training on vehicles can lead to prolonged waiting times for model aggregation, further exacerbating potential model obsolescence.
@@ I-B Motivation and Overview of Our Methodology
This work is among the first in the literature to address multi-model training within VEC-HFL. In our problem of interest, vehicles serve as training clients, multiple edge servers (ESs) act as middle-layer aggregators, and a cloud server (CS) functions as the global aggregator. We first note that a key challenge in facilitating multi-model training is mitigating model obsolescence risks while enhancing the utilization of distributed and dynamic training data from vehicles. In particular, we observe that traditional model aggregation methods in FL and HFL are mainly either synchronous or asynchronous. Synchronous aggregation waits for all clients to upload their models to maintain consistency, but this can introduce significant delays and increase the risk of model obsolescence. Conversely, asynchronous aggregation initiates the aggregation process as soon as a subset of clients uploads their models, effectively reducing latency. However, this method may suffer from model instability and struggle to fully utilize data from all participating clients, potentially impacting the convergence speed of the global model. To effectively utilizing all hierarchical data during training process, we introduce a hybrid synchronous-asynchronous aggregation rule for VEC-HFL. This rule operates on a per-task basis: at the edge layer, each ES performs an aggregation only after receiving all local models of a specific task from vehicles (synchronous approach), while at the cloud layer, the CS conducts a global aggregation after obtaining a subset of the corresponding ES models (asynchronous approach). This hybrid approach leverages the advantages of both synchronous and asynchronous methods. By synchronously aggregating at the edge layer, we ensure model consistency and fully utilize the local training data from all participating vehicles, enhancing model accuracy and convergence speed. Simultaneously, the asynchronous aggregation at the cloud layer reduces overall latency by not waiting for all ESs, which can have different delays of model aggregation based on the distribution of vehicles covered by them and their channel conditions, to report their models, thereby mitigating delays and the risk of model obsolescence. This combination provides a natural solution that balances the trade-offs between timeliness and model quality, making it particularly suitable for dynamic IoV environments where both factors are critical. Table I presents a summary of related studies, highlighting the key differences of our paper.
@@
Building on this hybrid aggregation rule, we then target minimizing the global time/latency costs while ensuring balanced training across diverse ML tasks over resource-limited vehicles. We formulate this problem as a min-max optimization that turns out to be NP-hard, making the obtaining of its solution challenging in dynamic IoV systems. To tackle this problem, we propose a stagewise solution called the Hybrid Evolutionary And gReedy allocaTion method (HEART). In Stage 1 of HEART, we optimize the assignment between model training tasks and heterogeneous resource-limited moving vehicles. To this end, leveraging the strong global search capabilities and computational efficiency of Particle Swarm Optimization (PSO) and Genetic Algorithms (GA), we develop an efficient combination of improved PSO and GA to achieve or closely approximate the optimal and balanced task assignment solution. In Stage 2 of HEART, we coordinate the training sequence/rank of assigned tasks on each selected vehicle to accelerate decision-making by designing a greedy-based task ranking optimization method.
TABLE I: A summary of related works
Related studies Environmental attributes Task count Scene attributes Training sequence
Stable Dynamic Single Multiple IoT VEC If optimized
[11][12][13][18] $\checkmark$ $\checkmark$ $\checkmark$
[14][15][16][19][22][23] $\checkmark$ $\checkmark$ $\checkmark$
[17][20][24] $\checkmark$ $\checkmark$ $\checkmark$
[8][25][28][30][32] $\checkmark$ $\checkmark$ $\checkmark$
Our Paper $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ I-C Summary of Contributions
$\bullet$ Hybrid Synchronous-Asynchronous Aggregation Rule for Multi-Model Training in VEC-HFL: We introduce an aggregation rule tailored for VEC-HFL. This hybrid approach performs synchronous aggregation at the ES and asynchronous aggregation at the CS. By doing so, it effectively leverages the rich data generated by vehicles while significantly mitigating the adverse effects of model obsolescence, thereby enhancing the overall timeliness of the training process across ML tasks.
@@ II Background and Preliminaries
The VEC-HFL architecture of our interest (see Fig. 1) comprises one CS as the global aggregator, multiple ESs as middle-layer aggregators, and multiple vehicles as model training clients moving between ESs. We collect the ESs via the set $\mathcal{M}=\{1,...,M\}$, where each ES $m\in\mathcal{M}$ has a certain communication coverage. Also, we denote the set of vehicles in the IoV system by $\mathcal{N}=\{1,...,N\}$, where each vehicle $n\in\mathcal{N}$ is assumed to be located inside the considered IoV region during the training process, moving between the ESs. Moreover, we denote the set of all ML tasks in the system by $\mathcal{J}=\{1,...,J\}$, which are expected to be scheduled and trained/executed over moving vehicles.
@@
where $\|\cdot\|$ represents the Euclidean 2-norm, and $\beta^{\left\langle j\right\rangle}$ represents a small positive value used to control the convergence criterion. Alg. 1 provides an overview of our designed hybrid synchronous-asynchronous aggregation rule for VEC-HFL.
@@ III-A Modeling the Time Overhead of Local and Edge Iterations
The non-task-training time primarily consists of (i) the local model upload time and (ii) any idle period from the completion of one task’s local iteration to the start of the next [^1] We have disregarded the communication delay of the downlink between ESs-to-vehicles and CS-to-ESs, due to the strong broadcasting capabilities of ESs and CS – similar to [MatchingGame, luo2020hfel].. However, modeling this time needs careful considerations because, during the $g^{\text{th}}$ global iteration, vehicles immediately transmit their local models to the associated ES once they finish training, and thus two scenarios may arise:
@@
where $r_{m,n;[g,k]}^{V2E,\left\langle j\right\rangle}$ represents the vehicle-to-edge data rate of vehicle $n$ under the coverage of ES $m$ during the $k^{\text{th}}$ edge iteration, which is related to the distance between the vehicle and ES and the channel conditions [Platoon]. Besides, the size of the local model (in bits) is denoted by $I_{m,n}^{V2E,\left\langle j\right\rangle}$.
Let $t^{inact,\langle j\rangle}_{m,n;[g,k]}$ denote the inactive time of the $n^{\text{th}}$ vehicle covered by ES $m$ after completing its local iteration at edge iteration $k$ of global aggregation $g$. It can be observed that this inactive time is a function of the completion time of edge iteration k1k-1 within global aggregation $g$ for a task $j$, which we denote it by $t_{m,n;[g,k-1]}^{end,\langle j\rangle}$. Mathematically, we obtain this inactive time as follows:
@@
$t_{m,n;[g,k]}^{end,\langle j\rangle}=\hskip-11.38109pt\sum_{j^{\prime}:~\mathcal{I}(\lambda_{m,n;[g]}^{\langle j^{\prime}\rangle})\leq\mathcal{I}(\lambda_{m,n;[g]}^{\langle j\rangle})}\hskip-11.38109ptt_{m,n;[g,k]}^{\langle j^{\prime}\rangle}\overline{x}_{m,n;[g]}^{\left\langle j^{\prime}\right\rangle},j\in\mathcal{J},$ (12)
@@ IV-A Key Constraint Design
$\hskip-7.96674pt\frac{N}{J}-\xi_{1}\leq\sum_{m\in\mathcal{M}_{g}^{\left\langle j\right\rangle}}\sum_{n\in\mathcal{N}_{m;[g,k]}^{\left\langle j\right\rangle}}\hskip-5.69054pt{\overline{x}_{m,n;[g]}^{\left\langle j\right\rangle}}\leq\frac{N}{J}+\xi_{2},\forall j\in\mathcal{J}.$ (17)
@@ IV-C Proposed Stagewise Scheduling through Hybrid Evolutionary And gReedy allocaTion (HEART)
Due to the integer/discrete nature of the optimization variables, $\bm{\mathcal{P}}$ belongs to a sub-category of integer programming (IP) problems which are known to be NP-hard. For example, except for $\overline{x}_{m,n;[g]}^{\left\langle j\right\rangle}$ that denotes a 0-1 integer, the variable $\lambda_{m,n;[g,k]}$ representing the training sequence across diverse ML tasks greatly complicates the problem. Moreover, the hybrid synchronous-asynchronous aggregation rule, and the constraints related to time limitations brought by vehicular mobility and the consideration on balanced training among different tasks further bring challenges to solution design. Accordingly, conventional algorithms such as branch and bound, as well as dynamic programming can lose their effectiveness. Also, after comprehensive investigations and tests, we found that the popular reinforcement learning (RL)-based methods can bring challenges to the performance of convergence, especially when having an increasing number of ML tasks and moving vehicles. Therefore, to tackle this problem, we propose a stagewise scheduling method called HEART, designed to achieve a near-optimal solution while achieving a timely manner.
@@
Stage 1: Task Scheduling via a Hybrid Evolutionary Approach
Heuristic algorithms such as PSO and GA are widely used due to their exceptional performance in scheduling and resource allocation problems [OffloadedT, DynamicSR, Swarm]. However, the complexity inherent in these problems, particularly the one considered in Stage 1 of this paper, often causes these algorithms to get trapped in local optima when they are used in isolation. To overcome this challenge, we propose a hybrid optimization algorithm that leverages the complementary strengths of both PSO and GA. In conventional PSO, particles tend to cluster prematurely, which weakens the algorithm’s global search capability [STATCOM, PSOAA]. Meanwhile, GA requires careful parameter tuning and frequently suffers from premature convergence [MADDPG, Assessment]. By integrating these two methods, we strive to achieve solutions that are globally optimal—or at least near-optimal—and avoid the pitfalls commonly associated with using a single method.
@@
S1.1 Overview and sketch of the PSO process: In our proposed task scheduling framework, a conventional PSO algorithm would search for optimal scheduling solutions across all vehicles at once [DynamicSR, STATCOM]. In particular, conventional PSO, inspired by the social behavior of bird flocks or fish schools, operates by having particles (representing solutions) move through the solution space. Each particle adjusts its position based on its own experience and that of its neighbors, guided by a fitness function that evaluates solution quality. The algorithm iteratively refines particle positions to converge on an optimal or near-optimal solution. While this global approach captures overall system objectives, it significantly increases the algorithm’s complexity and makes it difficult to tailor the scheduling solution for each individual vehicle in our setting of interest. Conversely, assigning tasks on a per-vehicle basis risks neglecting the overarching need for balanced global task allocation. A particular challenge arises because assigning more tasks can provide richer training data—beneficial for building a high performance global model—but since different tasks require different training times, within a limited dwell time, vehicles may often end up choosing tasks with shorter training durations to fit in as many tasks as possible when task assignments are made on a per-vehicle basis. As a result, tasks that take longer to train will be assigned less frequently, slowing down the progression of the global model accuracy and increasing thr total training time. In the following, we detail our proposed enhanced PSO.
@@
$f_{m,n;[g]}=\begin{cases}-\infty&\hskip-62.59605pt\text{if it can not meet (16)}\\ \sum_{j=1}^{J}\overline{x}_{m,n;[g]}^{\left\langle j\right\rangle}-(\xi_{3}\sum_{j=1}^{J}|\psi_{j}\\ ~~-\chi|-\sum_{j=1}^{J}\rho_{j}|\psi_{j}-\chi|)&\hskip-11.38109pt\text{otherwise}\end{cases}.$ (19)
@@
$\overline{x}_{m,n;[g]}^{p,\left\langle j\right\rangle}(\tau)$ $=\begin{cases}1&\text{if}~~e_{1}\leq\Phi^{p,\left\langle j\right\rangle}_{m,n;[g]}(\tau)\\ 0&\text{otherwise}\end{cases},$ (23)
@@
$\bullet$ Crossover enhancement. During the task scheduling phase for each vehicle, we sequentially select two adjacent particles, $p$ and $p+1$, correspond to task scheduling $\overline{x}_{m,n;[g]}^{{p,\left\langle j\right\rangle}}(\tau)$, $\overline{x}_{m,n;[g]}^{{p+1,\left\langle j\right\rangle}}(\tau)$. We then apply crossover and mutation on these allocations repeatedly. Specifically, for crossover, $p$ and $p+1$ serve as parent solutions, and we define a genetic length $J$ along with a randomly chosen crossover point $r$ such that $1<r<J$. Exchanging the segments of the parent allocations at the crossover point creates two offspring, $c_{1}$ and $c_{2}$. Thus, new task scheduling schemes are generated, which expand the exploration space and reduce the likelihood of trapping the algorithm in suboptimal solutions. We define all task scheduling for $p$, $p+1$ as a vector $X^{p}_{m,n;[g]}(\tau)$ and $X^{p+1}_{m,n;[g]}(\tau)$ (e.g., for $p$ during the $\tau^{\text{th}}$ iteration, if the particle implies the scheduling of tasks 1, 3, and 4 out of the four available tasks in the system, we will have the vector $X^{p}_{m,n;[g]}(\tau)=[1,0,1,1]$). To better describe this crossover process, we use $[i:j]$ to represent the elements from $i$ to j1j-1 in the vector, e.g., considering $X^{p}_{m,n;[g]}(\tau)=[1,0,1,1]$, we have $X^{p}_{m,n;[g]}(\tau)[0:2]=1,0$. We then obtain the offspring inherit allocations from the parents according to the following crossover rules: